OpenStudy (anonymous):

Kim has a strong first serve; whenever it is good (that is, in) she wins the point 70% of the time. Whenever her second serve is good, she wins the point 60% of the time. Seventy percent (75%) of her first serves and 85% of her second serves are good. What is the probability that Kim wins the point when she serves?

OpenStudy (anonymous):

if it is confusing ( and it looks confusing to me) try it with numbers lets say she serves \(1000\) times, and \(75\%\) are good. then she has \(.75\times 1000=750\) good serves, and of those \(70\%\) she wins, so she wins \(.7\times 750=525\) if the first serve is good now we consider what happens if the first serve is not good

OpenStudy (anonymous):

i am not a tennis player, but i assume that she only gets a second serve if her first serve is bad, which would be \(25\%\) of the time, and \(.25\times 1000=250\) of those \(250\) \(85\%\) are good, so \(.85\times 250=212.5\) second serves are good of those, \(60\%\) of the time she wins, making \(.6\times 212.5=127.5\) wins

OpenStudy (anonymous):

her total wins is therefore \(525+127.5=652.5\) take that out of \(1000\) to get a probability of \[.6525\]

OpenStudy (anonymous):

if you want to repeat this without the \(1000\) then compute \[.75\times .7+.25\times .85\times .6\]